1) Best Darn Glasses (BDR) is thinking of investing in a sandblasting machine for its glassware. It provides you with the following information: The initial investment for this project would be $235,000 in specialized machinery. According to CRA, this machine falls into a CCA class of 8%. There is the possibility of salvage of $6,000, although it’s not for sure. The risk-adjusted cost of capital is 12% and the company’s tax rate is 25%. Calculate the CCA tax shield under both scenarios – with and without salvage.

2) Using the information from above, calculate the project’s NPV if the following information were also provided to you: Cost of maintenance of the sandblasting machine is $35,000 per year, and the machine will only last 10 years. The salvage value, at that point, will be zero. The company’s revenues will be $170,000 per year with direct production costs of $27,000.

Please show all calculations and workings. If you are going to use Excel, please put both questions on the same sheet. You may also use Word. Please also name your file with your name before you submit it.
Answer #1
1…
EoYr. 0 1 2 3 4 5 6 7 8 9 10
1.Depn. At 20%(for Class 8)on carrying value 47000 37600 30080 24064 19251 15401 12321 9857 7885 6308
2.Carrying value of M/c 235000 188000 150400 120320 96256 77004.8 61603.84 49283.07 39426.46 31541.17 25232.93
3.Tax shield(Depn.in Row 1*25%) 11750 9400 7520 6016.00 4812.80 3850.24 3080.19 2464.15 1971.32 1577.06
4. PV F at 12%(1/1.12^Yr.n) 1 0.89286 0.79719 0.71178 0.63552 0.56743 0.50663 0.45235 0.40388 0.36061 0.32197
5. PV at 12%(4*5) 0 10491.07 7493.62 5352.59 3823.28 2730.91 1950.65 1393.32 995.23 710.88 507.77
6. NPV of CCAtax shields(Sum of row 5) 35449.32
(As the m/c has not reached the salvage of $ 6000 , the above workings are the same under both the scenarios-- with and without salvage
2..
Year 0 1
1.Initial Investment -235000
2.Revenues 170000
3.Direct prodn. Costs -27000
4.Maintenance -35000
5.EBT(sum 2 to 4) 108000
6. Tax at 25%(5*25%) -27000
7. EAT (5-6) 81000
Yr. 0 initial costs -235000
PV of EAT as above
81000*5.65022= 457668
(P/A,i=12%,n=10)
PV of tax shields(from 1 above) 35449
NPV of the purchase 258117

1.

Depn. In Yr. 1 for class 8 m/c=20%*235000= 47000
CCA tax shield in Yr. 1=47000*25%= 11750
Unlike under the straight-line method,under the declining balance method of calculating depreciation, salvage value , is not subtracted from the original cost , to know the depreciable base of the asset.
Under Decl.bal. like CCA, the cost is the depreciable base.
so, individually,both the depreciation amt. & the CCA tax shield amt. remain the same as above, with & without salvage.